Vapour pressure of aq solution of cane sugar is 75.6 cm of hg at 100c what is molality of solution.

Grade:10

1 Answers

Arun
6 years ago
The vapor pressure of pure water at 100°C is 76.0 cmHg. 
The mole fraction of water in the solution is: 
(75.6 cmHg) / (76.0 cmHg) = 0.994737 
The mole fraction of sugar in the solution is: 
1 - 0.994737 = 0.005263 
Take a hypothetical sample of exactly 1 mole of the solution: 
(0.994737 mol H2O) x (18.01532 g H2O/mol) = 17.9205 g = 0.0179205 kg 
(0.005263 mol sugar) / (0.0179205 kg) = 0.29368 mol/kg = 0.294 m sugar